# Smooth Tests for Normality in ANOVA

### v3

@@ Abstract
The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is seldom subjected to formal testing in practice. In this paper, we develop Neyman’s smooth tests for assessing normality in a broad class of ANOVA models. The proposed test statistics are constructed via the Gaussian probability integral transformation of ANOVA residuals and are shown to follow an asymptotic Chi-square distribution under the null hypothesis, with degrees of freedom determined by the dimension of the smooth model. We further propose a data-driven selection of the model dimension based on a modified Schwarz’s criterion. Monte Carlo simulations demonstrate that the tests maintain the nominal size and achieve high power against a wide range of alternatives. Our framework thus provides a systematic and effective tool for formally validating the normality assumption in ANOVA models.
Keywords: ANOVA; Estimation effect; Normality; Schwarz’s selection rule; Smooth test.
@@ 1 Introduction
Analysis of Variance (ANOVA) is a fundamental and widely used tool in both exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs. The theory of ANOVA has been well established in the literature; see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), and Montgomery, (2017). A standard assumption in ANOVA is that the random errors are normally distributed. Violations of this assumption may invalidate normal-based inference. From an estimation perspective, departures from normality undermine the validity of variance component estimators, since the classical variance formulae rely on mean squares of random effects being constant multiples of Chi-square variables (Scheffé,, 1959). From a testing perspective, the $F$-test in ANOVA is sensitive to nonnormality, especially when group sizes are unequal (Ali and Sharma,, 1996). The impact of nonnormality on the size and power of the $F$-test has been extensively studied (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964; Donaldson,, 1968; Tiku,, 1971; Janic-Wróblewska and Ledwina,, 2000). These considerations highlight the importance of rigorously evaluating the normality assumption in ANOVA models.
@@
Although a variety of normality tests have been proposed for observed data and linear models, relatively little attention has been paid to systematically assessing normality in ANOVA. A straightforward diagnostic approach is to use the normal probability plots of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which are intuitive but lack theoretical justification. To our knowledge, formal hypothesis testing procedures designed explicitly for ANOVA are scarce in the literature. Most existing methods simply apply classical normality tests to ANOVA residuals, see, for example, Bonett and Woodward, (1990), Bonett and Seier, (2002), and Hwang and Wei, (2006). While applicable in practice, such approaches do not provide a rigorous assessment of normality within the ANOVA framework and fail to account for the structural constraints inherent in ANOVA models.
@@
In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing normality in various types of ANOVA models. Neyman’s smooth test is widely used in diverse scientific fields due to its theoretical soundness and practical effectiveness, with proven power against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022). Specifically, we reformulate the normality testing problem as one of uniformity via the Gaussian probability integral transform (PIT), and construct test statistics based on the Gaussian PIT of ANOVA residuals. Under mild conditions, the proposed statistics are asymptotically Chi-square distributed under the null hypothesis. We further analyze the power properties under both fixed and local alternatives. A modified Schwarz’s selection rule is proposed to determine the direction of tests. Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
@@
The main contributions of this paper are threefold. First, the proposed asymptotic Chi-square tests are simple to implement: critical values are obtained directly from the limiting distributions without resorting to resampling. Second, we systematically investigate the properties of the test statistics in three types of one-way fixed effects models, potentially with a diverging number of groups, within a unified asymptotic framework. Our analysis clarifies how parameter estimation affects the test statistics: heterogeneity in group means and variances alters their explicit form and the corresponding regularity conditions, but the convergence rate depends only on the total sample size under both null and alternative hypotheses. Third, we provide theoretical justification for data-driven smooth tests in the ANOVA setting, including a revised limiting null distribution for the corresponding statistics in finite samples.
@@
The remainder of this paper is organized as follows. Section 2 introduces the general testing framework. Section 3 establishes theoretical results for three types of ANOVA models. Section 4 proposes a data-driven testing procedure. Section 5 reports simulation results. Section 6 discusses further extensions. Section 7 concludes. Proofs and additional numerical results are provided in Appendix A.
@@ 2 The testing framework
Our problem of interest is to determine whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$, that is, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ for some $\sigma>0$. Equivalently, this can be formulated as testing the null hypothesis about the standardized random error $H_{0}:e\sim\mathcal{N}(0,1)$ against alternatives of nonnormality. To motivate our methodology, consider the transformation $Z=\Phi(e)$, where $\Phi(\cdot)$ denotes the cumulative distribution function (CDF) of the standard normal distribution. Under the null hypothesis of normality, the CDF of the transformed random variable $Z$ is given by
$G(z)=\mathrm{P}(Z\leq z)=\mathrm{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1,z\in[0,1]$, that is, $Z\sim\mathcal{U}[0,1]$, where $\mathcal{U}[0,1]$ denotes the uniform distribution on the interval $[0,1]$. Under the alternative, the density $g(z)$ deviates from unity. This distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the foundation for Neyman’s smooth test. In particular, Neyman, (1937) introduced the following smooth alternative to the uniform density:
@@
where $c(\bm{\theta}_{K})$ is a a normalizing constant depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ denotes an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
@@
If independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ are available and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
@@
However, in practice, the sequence $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown. We instead rely on approximations $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and an estimator $\hat{\sigma}$ for $\sigma$ (e.g., residuals and the sample standard deviation, respectively). Accordingly, we consider the following approximation for $Z_{i}$:
$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n,$
@@
$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$ (2.4)
The random variables $\{\hat{Z}_{i}\}_{i=1}^{n}$ are not i.i.d. any longer due to estimation effects caused by $\hat{\varepsilon}_{i}$ and $\hat{\sigma}$. Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$. Compared with $\Psi_{K}^{2}$ in (2.3), the statistic based on $\hat{Z}_{i}$ suffers from non-negligible estimation effects, as demonstrated in the theoretical results below. Consequently, the presence of estimation effects invalidates the form of the “naive” statistic (2.4) and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution. This motivates a detailed analysis of estimation effects, leading to the correct quadratic-form test statistic.
Another important issue concerns sample sizes in ANOVA models. To highlight the essence of smooth tests, we first focus on the simplified setting of a single random sequence with sample size $n$ throughout this section. This simplified setting captures the core methodological ideas and facilitates the derivation of key asymptotic properties. In contrast, the complete ANOVA setting involves additional complications arising from multiple sample sizes and the associated indexing structure, which demand more refined analysis. From a practical standpoint, our method is further designed to accommodate scenarios where the number of groups increases with the total sample size. These issues will be systematically investigated in subsequent analysis of specific ANOVA models.
@@ 3 Testing for normality in one-way fixed effects models
In this section, we develop smooth tests for normality in three types of one-way fixed effects models, where each model assumes equality in at least one of the group means or the group variances. The testing procedure and its theoretical properties are investigated separately for each case. Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$. The total sample size is defined as $N=\sum_{j=1}^{J}N_{j}$. In our theoretical framework, the number of groups, $J$, is allowed to diverge as $N$ increases.
3.1 Same mean and same variance
We begin with the ANOVA models where both the group means and group variances are equal. Specifically, consider the model
@@ 3.1 Same mean and same variance
The parameters $\mu$ and $\sigma^{2}$ in (3.1) are estimated by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To derive the feasible test statistic and its properties, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ for $k=1,\ldots,K$. For this purpose, we impose the following two assumptions.
@@
The two mild assumptions are in line with similar ones adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022). For subsequent analysis, we introduce two constants:
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is equivalent to the sum of three terms after neglecting the higher-order term: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$. The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$ (3.2)
$\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
@@
$\sigma_{kl}$ $=\mathrm{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
@@
$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.3)
The following corollary establishes the limiting null distribution of $\hat{\Psi}_{K}^{2}$ defined in (3.3).
@@
Corollary 1 establishes an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$. Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
$N\hat{\Psi}_{K}^{2}>\chi_{K,1-\alpha}^{2},$
@@
We now investigate the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives as well as under a Pitman-type sequence of local alternatives. For the fixed alternatives, we consider the following form:
@@
Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed. Consequently, the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ differs from that under $H_{0}$. To formally state the corresponding theoretical results under $H_{1}$, we introduce the following notations. Let
$d_{1k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)f(x)\mathrm{d}x,$
$d_{2k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)xf(x)\mathrm{d}x,$
$d_{3k}=\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)xf(x)\mathrm{d}x,$
@@
$d_{4k}=\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
for $k=1,\ldots,K$. Note that under $H_{0}$, the constants $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ since $f(x)=\phi(x)$. The following theorem characterizes the asymptotic properties of $\hat{\Psi}_{K}^{2}$ under the alternative hypothesis $H_{1}$.

### v4

# Data-driven Smooth Tests for Normality in ANOVA When the Number of Groups is Large

@@ Abstract
The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models. However, it is rarely subjected to formal testing in practice, and theoretically justified procedures are largely unavailable, especially when the number of groups diverges. In this paper, we develop Neyman’s smooth tests for assessing normality in a broad class of ANOVA models, allowing the number of groups to diverge. The proposed test statistics are constructed via the Gaussian probability integral transformation of ANOVA residuals. We show that using residuals induces non-negligible parameter estimation effects, whose structure depends on the underlying ANOVA model and plays a crucial role in shaping the form of the test statistics and their asymptotic behavior. Under the null hypothesis of normality, the resulting statistics follow an asymptotic Chi-square distribution, with degrees of freedom determined by the order of the smooth test (i.e., the number of components included in the smooth test). We further propose a modified Schwarz’s selection rule to automatically determine the order, thereby yielding fully data-driven smooth tests that require no additional tuning parameters. Simulation studies and a real-data example indicate that the proposed tests perform well in practice and are readily applicable.
Keywords: ANOVA; Estimation effects; Normality; Schwarz’s selection rule; Smooth tests
@@ 1 Introduction
Analysis of Variance (ANOVA) is a fundamental and widely used tool in both exploratory and confirmatory data analysis (gelman2005analysis), particularly for comparing group means and assessing the significance of factors in experimental designs. The theory of ANOVA has been well established in the literature; see, for example, scheffe1959anova, miller1997beyond, wickens2004design, dean2017design, hirotsu2017advanced, and montgomery2017design. A standard assumption in ANOVA is that the random errors are normally distributed. Violations of this assumption may invalidate normal-based inference. From an estimation perspective, departures from normality undermine the validity of variance component estimators, since the classical variance formulae rely on mean squares of random effects being constant multiples of Chi-square variables (scheffe1959anova). From a testing perspective, the $F$-test in ANOVA is sensitive to nonnormality, especially when group sample sizes are unbalanced (ali1996robustness) or the number of groups is large (akritas2000asymptotics; akritas2004heteroscedastic; wang2006two). The impact of nonnormality on the size and power of the $F$-test has been extensively studied (pearson1931analysis; gayen1950distribution; david1951method; david1951effect; srivastava1959effect; atiqullah1962estimation; tiku1964approximating; donaldson1968robustness; tiku1971power; akritas2000asymptotics; janic2000data). These considerations highlight the importance of rigorously evaluating the normality assumption in ANOVA models.
@@
Although a variety of normality tests have been proposed for observed data and linear models (see, for example, dagostino1986goodness and bonett2002test), relatively little attention has been paid to the systematic assessment of normality in ANOVA. A straightforward diagnostic approach is to use the normal probability plots of ANOVA residuals (dean2017design; hirotsu2017advanced), which is intuitive but lacks theoretical justification. To our knowledge, formal normality testing procedures explicitly designed for ANOVA are scarce in the literature. Most existing methods instead apply classical normality tests—such as the Shapiro–Wilk test (shapiro1965analysis; shapiro1972approximate), the Jarque–Bera test (jarque1987test), and the Kolmogorov–Smirnov test (in particular, the Lilliefors version) (lilliefors1967kolmogorov)—directly to ANOVA residuals; see, for example, bonett1990testing; bonett2002test; hwang2006novel. While convenient in practice, such approaches suffer from two fundamental limitations. First, they ignore the estimation effect induced by fitting the ANOVA model, which may alter the distributional properties of the residuals. Second, they lack theoretical guarantees, especially in ANOVA models with a diverging number of groups, under which these procedures may become invalid. Consequently, they do not provide a rigorous assessment of normality within the ANOVA framework and fail to account for the inherent structural constraints of ANOVA.
@@
In this article, following the spirit of neyman1937smooth, we propose a unified Neyman’s smooth test framework for assessing the normality of random errors in various types of ANOVA models where the number of groups is allowed to diverge. Neyman’s smooth tests are widely used in diverse scientific fields due to their theoretical soundness and practical effectiveness; see, e.g., bera_neymans_2002, Chapters 4 and 10 of thasComparingDistributions2010, and Section 16.4 of Lehmann_Romano_2022. Specifically, we reformulate the normality testing problem as a uniformity problem via the Gaussian probability integral transform (PIT) and construct test statistics based on the Gaussian PIT of the ANOVA residuals. The use of residuals introduces parameter estimation effects whose structure depends on the underlying model and is therefore intrinsic to the ANOVA framework. Consequently, the resulting statistic takes the form of a quadratic expression involving the inverse of a covariance matrix, whose structure is determined by these estimation effects. Given mild conditions, the proposed test statistics are asymptotically Chi-square distributed under the null hypothesis of normality. We also analyze the power properties under both fixed and local alternatives. Furthermore, a modified Schwarz’s selection rule is proposed to determine the test directions (i.e., the order of the test), yielding fully data-driven smooth tests that do not require additional tuning parameters. Numerical results demonstrate the good performance of our proposed methodology in finite samples.
@@
The main contributions of this paper are twofold. First, we systematically investigate the effects of parameter estimation on ANOVA residuals-based normality testing. We show that heterogeneity in group means and variances alters the explicit form of the test statistics and the associated regularity conditions, while the convergence rate depends solely on the total sample size under both the null and alternative hypotheses. Second, we provide rigorous theoretical justification for data-driven smooth tests in the ANOVA setting. In particular, we establish a revised limiting null distribution for the corresponding statistics that remains valid in finite samples. As a result, the proposed tests are simple to implement, with critical values obtained directly from the limiting distributions without resorting to resampling. Overall, the proposed framework accommodates settings with a potentially diverging number of groups and thus overcomes key limitations of many classical methods.
@@
The remainder of this paper is organized as follows. Section 2 introduces the general testing framework. Section 3 establishes theoretical results for three types of ANOVA models. Section 4 proposes a data-driven testing procedure. Section 5 reports simulation results. Section 6 provides an empirical application. Section 7 concludes. Proofs and additional numerical results are provided in the Online Appendix.
@@ 2 The testing framework
Our problem of interest is to assess whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$, that is, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ for some $\sigma>0$. Equivalently, this can be formulated as testing the null hypothesis about the standardized random error $H_{0}:e\sim\mathcal{N}(0,1)$ against alternatives of nonnormality. To motivate our methodology, consider the transformation $Z=\Phi(e)$, where $\Phi(\cdot)$ denotes the cumulative distribution function (CDF) of the standard normal distribution. Under the null hypothesis of normality, the CDF of the transformed random variable $Z$ is given by
$G(z)=\mathbb{P}(Z\leq z)=\mathbb{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1,z\in[0,1]$, that is, $Z\sim\mathcal{U}[0,1]$, where $\mathcal{U}[0,1]$ denotes the uniform distribution on the interval $[0,1]$. Under the alternative, the density $g(z)$ deviates from unity. This distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ forms the basis of Neyman’s smooth test. In particular, neyman1937smooth introduced the following smooth alternative to the uniform density:
@@
where $c(\bm{\theta}_{K})$ is a normalizing constant depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ denotes an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
@@
If independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ are available and $\sigma$ is also known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
@@
However, in practice, the sequence $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown. We instead rely on approximations $\{\widehat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and an estimator $\widehat{\sigma}$ for $\sigma$ (e.g., residuals and the sample standard deviation, respectively). Accordingly, we consider the following approximation for $Z_{i}$:
$\widehat{Z}_{i}=\Phi(\widehat{e}_{i})=\Phi\left(\frac{\widehat{\varepsilon}_{i}}{\widehat{\sigma}}\right),\quad i=1,\ldots,n,$
@@
$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\widehat{Z}_{i}\right)\right)^{2}.$ (2.4)
The random variables $\{\widehat{Z}_{i}\}_{i=1}^{n}$ are not i.i.d. any longer due to estimation effects caused by $\widehat{\varepsilon}_{i}$ and $\widehat{\sigma}$. Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\widehat{Z}_{i})$. Compared with $\Psi_{K}^{2}$ in (2.3), the statistic based on $\widehat{Z}_{i}$ suffers from non-negligible estimation effects, as demonstrated in the theoretical results below. Consequently, the presence of estimation effects invalidates the form of the “naive” statistic (2.4) and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution. This motivates a detailed analysis of the effects of estimation that yields the correct quadratic-form test statistic.
Another important issue concerns sample sizes in ANOVA models. To highlight the essence of smooth tests, we first focus on the simplified setting of a single random sequence with sample size $n$ throughout this section. This simplified setting captures the core methodological ideas and facilitates the derivation of key asymptotic properties. In contrast, the complete ANOVA setting introduces additional complications due to multiple sample sizes and the associated indexing structure, necessitating a more refined analysis. From a practical standpoint, our method is designed to accommodate scenarios in which the number of groups increases with the total sample size. These issues will be systematically investigated in subsequent analysis of specific ANOVA models.
@@ 3 Testing for normality in one-way fixed effects models
In this section, we develop smooth tests for normality in three types of one-way fixed effects models, each imposing homogeneity on either the group means or the group variances. The testing procedures and their theoretical properties are studied separately for each case. Formally, let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observed responses from $J$ groups, where $N_{j}$ is the sample size of group $j$ for $1\leq j\leq J$. The total sample size is $N=\sum_{j=1}^{J}N_{j}$. In our asymptotic framework, the number of groups $J$ is allowed to diverge as $N\to\infty$.
3.1 Common group means and common group variances
We begin with ANOVA models in which both the group means and group variances are equal. Specifically, consider the model
@@ 3.1 Common group means and common group variances
The parameters $\mu$ and $\sigma^{2}$ in model (3.1) are estimated by the sample mean $\widehat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\widehat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\widehat{\mu}\right)^{2}$, respectively. Define $\widehat{\varepsilon}_{ij}=Y_{ij}-\widehat{\mu}$, $\widehat{e}_{ij}=\widehat{\varepsilon}_{ij}/\widehat{\sigma}=(Y_{ij}-\widehat{\mu})/\widehat{\sigma}$, and $\widehat{Z}_{ij}=\Phi(\widehat{e}_{ij})$. To derive the feasible test statistic and its properties, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ for $k=1,\ldots,K$. For this purpose, we impose the following two assumptions.
@@
The two mild assumptions are in line with similar ones adopted in the literature of Neyman’s smooth tests, including bera2013smooth and song2022smooth. For subsequent analysis, we introduce two constants:
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ is equivalent to the sum of three terms after neglecting the higher-order ones: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$. The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$, $k=1,\ldots,K$.
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)\right]^{\top}$ (3.2)
$\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
@@
$\sigma_{kl}$ $=\mathbb{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
@@
$\widehat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right).$ (3.3)
The following corollary establishes the limiting null distribution of $\widehat{\Psi}_{K}^{2}$ defined in (3.3).
@@
Corollary 1 establishes an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\widehat{\Psi}_{K}^{2}$. Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
$N\widehat{\Psi}_{K}^{2}>\chi_{K,1-\alpha}^{2},$
@@
We now investigate the asymptotic behavior of $\widehat{\Psi}_{K}^{2}$ under the fixed alternatives as well as under a Pitman-type sequence of local alternatives. For the fixed alternatives, we consider the following form:
@@
Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed. Consequently, the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ and $\widehat{\Psi}_{K}^{2}$ differs from that under $H_{0}$. To formally state the corresponding theoretical results under $H_{1}$, we introduce the following notations. Let
$d_{1k}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)f(x)\mathrm{d}x,$
$d_{2k}=\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)xf(x)\mathrm{d}x,$
$d_{3k}=\mathbb{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)xf(x)\mathrm{d}x,$
@@
$d_{4k}=\mathbb{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
for $k=1,\ldots,K$, where $\phi$ denotes the PDF of standard normal distribution. Note that under $H_{0}$, the constants $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ since $f(x)=\phi(x)$. The following theorem characterizes the asymptotic properties of $\widehat{\Psi}_{K}^{2}$ under the alternative hypothesis $H_{1}$.